1. Determine the electric length of a 5cm long Wifi 2.4 GHz antenna, in degrees.

2. VSWR of a system defined as ZL = (1+j2) Z0

3. Find the input impedance of a 50 Ohm quarter wavelength long TL is terminated with a load of 100 Ohms.
Answer #1

» 20 0 Sleef 2 a= c/f = 3x10 x 10² cm Qlochic length Electric length = [Pl = 2y x se= longth] f = c/ 2.48109 2 = 12.5cm Subst3) given e=My 20 - 50 2 = 100 when lean input impedance forom wla is Zijp = 20%. = 50x50 100 zipp = 250]